In the text, we calculated the change in real GDP in the hypothetical economy of Table 2-3. using the prices of 2005. Calculate the change in real GDP between 2005 and 2010 using the same data but the prices of 2010. Your answer should demonstrate that the prices that are used to calculate real GDP do affect the calculated growth rate, but typically not by very much.



1
Expert's answer
2021-08-09T06:49:26-0400


Nominal GDP is the GDP calculated on the current year prices and the real GDP is the GDP calculated on the prices of base year. The base year is chosen to nullify the effect of price rise on the calculation of GDP. Real GDP gives a better picture of the economy’s growth. Nominal GDP gives the inflated picture as it is affected by the inflation.

The data given:



Step I: The nominal GDP (NGDP) for year 2012 would be output produces in an economy in year 2012 multiplied by prices in the current year i.e. 2012.

"NGDP_{2012}=GDP_{2012} \times Price_{2012}"

Step II: The real GDP for year 2012 would be output produces in an economy in year 2012 multiplied by prices in the base year. In case we take i.e. 2000 as a base year, the real GDP (RGDP) is

"RGDP_{2012}=GDP_{2012} \times Price_{2000}"

Step III: Change in real GDP (RGDP) between 2000 and 2012 is the difference in real GDP of current (2012) and base year (2000), divided by real GDP of base year (2000).

"ΔRGDP= \frac{(Real \; GDP_{2012}- Real \; GDP_{2000})}{Real \;GDP_{2000}}"

Step IV: The change in real GDP with 2000 prices can be calculated as follows:

"ΔRGDP= \frac{3.50-1.50}{1.50} \times 100 \\

= 1.33 \times 100 \\

= 133 \; \%"

Step V: Now to calculate the real GDP on 2012 prices, we need to calculate the GDP of 2000 on 2012 prices:

We take quantity of year 2012 and prices of year 2000. Thus

For 1 beer priced at $1.50

1 at $1.50=$1.50

For 1 skittles priced at $1.00

1 at $1.00 = $1.00

Total real GDP in 2000 on 2012 prices

Total=$2.50


Step VI: To calculate the change in real GDP of 2012 as 6.00 and real GDP of 2000 as 2.50

"ΔRGDP=\frac{6.00-2.50}{2.50} \times 100 \\

= 1.40 \times 100 \\

= 140 \; \%"

The change in real GDP with 2000 Prices was 133% and the change in real GDP with 2012 prices was 140%. Thus, the prices that are used to calculate real GDP does affect the calculated growth rate but the variance is not too much. The prices that are used to calculate real GDP do affect the calculated growth but not too much.